OpenStudy (anonymous):

A spring has a force constant of 80000 N/m. How far must it be stretched for its potential energy to be 21 J? Answer in units of m.

OpenStudy (mayankdevnani):

P.E=mgh F=mg

OpenStudy (mayankdevnani):

21=80000*h h=21/80000=0.0002625 m

OpenStudy (anonymous):

So, is the distance 0.0002625?